# On the Effects of Data Scale on Computer Control Agents

### v3

@@ 4.3 In-domain performance
We start by evaluating zero-shot, few-shot and LoRA-tuned methods in domain. Table 4 reports the step-wise accuracy performance on the IDD sub-split of Random-500. In-domain, LoRA-tuned models, despite using the smaller PaLM 2S model, when trained with sufficient amounts of data, largely outperform the zero-shot and few-shot methods. For low-level instructions, even LT-5 surpasses all non-fine-tuned models, while for high-level instructions, it requires more training data (1k episodes). The best fine-tuned model reaches 71.5% on high-level and 86.6% on low-level.
@@ 4.4 Effect of scale on in-domain transfer
(a) In-domain
(b) Out-of-domain
Figure 5: Relationship between step-wise accuracy and number of fine-tuning samples. With 5–100 episodes we use rank=4; with $\geq$1k episodes we use rank=64. IDD performance is based on the full IDD test split. OOD performance is based on the average across the 3 OOD splits.
Figure 5(a) visualizes the number of training episodes drawn from AndroidControl and the step accuracy achieved on the full IDD test sub-split. Both high and low-level curves exhibit linear trends with the log of training data. We extrapolate that it would take 500K and 1M episodes to reach $95\%$ step-accuracy for low and high-level instructions, respectively. However, while low-level tasks can be accomplished in one step, high-level tasks require multiple steps. Taking 5 steps as the rough length for high-level tasks, we predict 2M episodes would be required to reach the 99% step-wise accuracy to achieve 95% episode completion for high-level tasks (see Appendix E.4 for an empirical evaluation of performance as episode length is varied). While this is conservative by assuming no possible mistake recovery, we still feel this analysis provides a helpful rough quantification.
@@ 4.5 Effect of scale on out-of-domain transfer
Table 5: OOD step accuracy. In square brackets [X] we report the delta from the IDD accuracy obtained on the full IDD test split.
IDD app-unseen task-unseen cat-unseen
@@
As the number of fine-tuning samples increases, performance improves and so does the gap between IDD and OOD performance (Table 5). With 10k or more episodes, the IDD accuracy is noticeably higher than on the three ODD splits. For example, with LT-10k, the gap is 7.7–8.7 pp for high-level instructions and 2.0–3.2 pp for low-level instructions. In general, more out-of-domain transfer occurs for low-level tasks, which is expected as low-level tasks share more similarity across tasks and apps than high-level tasks.
As for in-domain, we extrapolate how much training data would be necessary to achieve a reasonable accuracy out of domain (Figure 5(b)). OOD step-accuracy grows more slowly than in domain, and is estimated to reach 95% at 10M and 60M episodes for low-level and high-level instructions respectively. Similar to above, the number of episodes we predict would be required to reach 99% step accuracy to therefore achieve 95% episode completion rate on 5-step high-level tasks is 150M. Based on these projections, it seems expensive but feasible to obtain good general LL performance with fine-tuning, while the predicted higher order of magnitude of the number of required demonstrations suggests fine-tuning alone may not be sufficient to achieve robust OOD performance on HL tasks.
@@ 5 Conclusion
6 Limitations
There are multiple potential limitations in this work. First, we only fine-tuned one model, PaLM-2S; however, while the absolute performance values would change, we expect our relative findings to be consistent across model families. Additionally, using offline evaluation for agent performance has the known issue of not rewarding alternative routes to complete a task and the ability to take corrective actions [3]. Finally, while selected to encompass important use cases, the set of app categories in AndroidControl is still an incomplete representation of all tasks users may ask agents to perform.
@@ Appendix A Ethical considerations
Autonomous computer control agents can bring value to visually-impaired users, by providing them with access to a much wider range of applications and functionality. More broadly, they can enhance human productivity by automating everyday tasks. Computer agents have societal, security and privacy implications. An agent may leak private information or carry out a task in an unacceptable way or produce unwanted side effects. Malicious actors could also use these agents for undesired purposes such as overriding anti-fraud mechanisms or manipulating applications to achieve undesirable goals. For these reasons, deployment of this technology going forward will have to be carefully considered and combined with research in other areas on LLM safety to balance potential societal trade-offs with risks.
@@ B.3 Accessibility node metadata
Each node in the Android accessibility tree corresponds to a UI element in the screen. Each node is described by multiple metadata. In our computer control agent implementation we use the following node metadata:
@@
Appendix C Computer control agent implementation
@@ C.2 Action space
type_text {"action_type":"type","text":<text_input>,"x":<x_coordinate>,"y":<y_coordinate>}
@@ D.1 Data processing and training details
To run our experiments we generate 2 SeqIO tasks (HL and LL) that process the AndroidControl dataset as follows. First, in order to support prediction of task completion actions not present in the original dataset, at the end of every episode we artificially insert a terminate action that takes the episode status (successful or infeasible) as argument. Second, we discard any step with an element-based action (click, long_press) that does not have a UI element associated. This is due to either a touch on an empty area or a target UI element missing from the accessibility tree. However, these discarded steps are still considered in the action history to support prediction of later steps that reference previous actions or elements. Finally, we discard (only from SeqIO LL tasks) steps that are missing a low-level instruction which annotators may have forgotten to enter. The discarded steps account for roughly 10% of all steps in the full dataset, while the equivalent number of episodes in the SeqIO tasks are not significantly affected as an episode is dropped only if all of its steps are discarded. Table 7 shows the statistics of the SeqIO tasks that are the results of this processing.
Table 7: Statistics of the SeqIO tasks generated from AndroidControl. Note that the number of steps here reported is higher than what reported in Table 3 because it accounts for terminate actions.
@@ D.2 LLM prompts for zero-shot experiments
For an LLM-based computer control agent, its prompt describes what the agent is expected to do, the action space and the expected format of actions, the task instruction, the current screen and application, and the history of previously executed actions and their outcome.
@@ D.3 Action matching
In computing step-wise accuracy, we consider an action correctly predicted if it matches the ground truth action exactly (i.e., action type and arguments are identical) or if it aligns with it as follows.
@@ E.2 Training with different levels of instructions
We observe the screen representations provided as input to the computer control agent are critical to its success. Table 13 compares the step accuracy of various models on Google’s first party apps and apps developed by third party developers (3rd-party). It is evident the gap in performance, especially in handling high-level instructions. All zero-shot methods perform significantly better on first-party apps than on third-party apps. This is evident for fine-tuned models, especially on high-level instructions where the most performance model (LT-all) achieves 82.5% step accuracy on first-party apps and only 58.7% accuracy on third-party apps. When the input is made of low-level instructions the gap is smaller and in some cases (few-shots) the performance on third-party apps is higher. This shows how for tasks that require stronger reasoning capabilities accurate screen representations are particularly critical.

### v5

@@ 4.3 In-domain performance
We start by evaluating zero-shot, few-shot and LoRA-tuned methods in domain. Table 4 reports the step-wise accuracy performance on the IDD sub-split of Random-500. In-domain, LoRA-tuned models, despite using the smaller PaLM 2S model, when trained with sufficient amounts of data, largely outperform the zero-shot and few-shot methods. For low-level instructions, even LT-5 surpasses all non-fine-tuned models, while for high-level instructions, it requires more training data (1k episodes). The best fine-tuned model reaches 71.5% on high-level instructions and 86.6% on low-level instructions.
@@ 4.4 Effect of scale on in-domain transfer
(a) In-domain. High-level trendline $R^{2}=0.999$, low-level trendline $R^{2}=0.951$.
(b) Out-of-domain. High-level trendline $R^{2}=0.986$, low-level trendline $R^{2}=0.987$.
Figure 5: Relationship between step-wise accuracy and number of fine-tuning samples. With 5–100 episodes we use LoRA rank=4; with $\geq$1k episodes we use LoRA rank=64. IDD performance is based on the full IDD test split. OOD performance is based on the average across the three OOD splits.
Figure 5(a) visualizes the number of training episodes drawn from AndroidControl and the step accuracy achieved on the full IDD test sub-split. Both high and low-level curves exhibit linear trends ($R^{2}$ coefficients are greater than 0.95) with the log of training data. We extrapolate that it would take 500K and 1M episodes to reach $95\%$ step-accuracy for low and high-level instructions, respectively. However, while low-level tasks can be accomplished in one step, high-level tasks require multiple steps. Taking 5 steps as the rough length for high-level tasks, we predict 2M episodes would be required to reach the 99% step-wise accuracy to achieve 95% episode completion for high-level tasks (see Appendix E.4 for an empirical evaluation of performance as episode length is varied). While this is conservative by assuming no possible mistake recovery, we still feel this analysis provides a helpful rough quantification.
@@ 4.5 Effect of scale on out-of-domain transfer
Table 5: OOD step accuracy. In square brackets [X] we report the percentage point increase (+) or drop (-) from the IDD accuracy obtained on the full IDD test split.
IDD app-unseen task-unseen categ-unseen
@@
As the number of fine-tuning samples increases, performance improves and so does the gap between IDD and OOD performance (Table 5). With 10k or more episodes, the IDD accuracy is noticeably higher than on the three ODD splits. For example, with LT-10k (r=4), the gap is 7.7–8.7 pp for high-level instructions and 2.0–3.2 pp for low-level instructions. With r=64 the gap increases. In general, more out-of-domain transfer occurs for low-level tasks, which is expected as low-level tasks share more similarity across tasks and apps than high-level tasks.
As for in-domain, we extrapolate how much training data would be necessary to achieve a reasonable accuracy out of domain (Figure 5(b)). OOD step-accuracy grows more slowly than in domain, and is estimated to reach 95% at 10M and 60M episodes for low-level and high-level instructions, respectively. Similar to above, the number of episodes we predict would be required to reach 99% step accuracy to therefore achieve 95% episode completion rate on 5-step high-level tasks is 150M. Based on these projections, it seems expensive but feasible to obtain good general LL performance with fine-tuning, while the predicted higher order of magnitude of the number of required demonstrations suggests fine-tuning alone may not be sufficient to achieve robust OOD performance on HL tasks.
@@
5 Limitations
There are multiple potential limitations in this work. First, we only fine-tuned one model, PaLM-2S; however, while the absolute performance values would change, we expect our relative findings to be consistent across model families. Additionally, using offline evaluation for agent performance has the known issue of not rewarding alternative routes to complete a task and the ability to take corrective actions [3, 27]. Further, while selected to encompass important use cases, the set of app categories in AndroidControl is still an incomplete representation of all tasks users may ask agents to perform. Finally, studying inference costs is not a focus of the paper, although it is obvious that it is much cheaper to predict on a fine-tuned PaLM-2S model than on any larger models, such as PaLM-2L or GPT-4.
@@ Appendix A Ethical considerations
Autonomous UI control agents can bring value to visually-impaired users, by providing them with access to a much wider range of applications and functionality. More broadly, they can enhance human productivity by automating everyday tasks. UI agents have societal, security and privacy implications. An agent may leak private information or carry out a task in an unacceptable way or produce unwanted side effects. Malicious actors could also use these agents for undesired purposes such as overriding anti-fraud mechanisms or manipulating applications to achieve undesirable goals. For these reasons, deployment of this technology going forward will have to be carefully considered and combined with research in other areas on LLM safety to balance potential societal trade-offs with risks.
@@ B.3 Accessibility node metadata
Each node in the Android accessibility tree corresponds to a UI element in the screen. Each node is described by multiple metadata. In our UI control agent implementation we use the following node metadata:
@@
B.4 Number of UI elements in test splits
In Table 3, we note that in the test sub-split that is sampled from the same distribution as the train split (the “IDD” split), the average number of UI elements per screen is comparable, as expected, while in the other 3 out-of-domain test sub-splits there is a significant drop in the number of UI elements per screen. To explain this phenomenon, we calculated the average number of UI elements across all screens in the data for each app (for simplicity we call this “the average number of UI elements per app”). Figure 7 shows the distribution of the average number of UI elements per app for the entire dataset. The distribution shows a long tail, with a few apps with a very large number of UI elements per screen. If we remove these outliers by dropping the top 5% of apps with the most elements per screen, the average number of UI elements across the remaining apps is 180.9, which is inline with the out-of-domain test sub-splits. A reasonable question is then why do the out-of-domain test sub-splits not sample from the tail? We believe this is because, as shown in Table 3, the number of apps in each sub-split is relatively small. Further, there is overlap between the different test sub-splits (this was done to keep the total size of the test set reasonable, if each sub-split needed to be disjoint, it would leave less train data), so the total number of apps across the union of all out-of-domain test sub-splits is still modest, again supporting that we simply did not draw enough outliner apps from the long tail to affect the average for the out-of-domain splits.
Figure 7: Histogram of average number of UI elements per app for all apps in the AndroidControl dataset. Please note the y-axis is logarithmic.
Appendix C UI control agent implementation
@@ C.2 Action space
type {"action_type":"type","text":<text_input>,"x":<x_coordinate>,"y":<y_coordinate>}
@@ D.1 Data processing and training details
To run our experiments we generate 2 SeqIO tasks (HL and LL) that process the AndroidControl dataset as follows. (i) In order to support prediction of task completion actions not present in the original dataset, at the end of every episode we artificially insert a terminate action that takes the episode status (successful or infeasible) as argument. (ii) The dataset omits identifiers for target elements associated with click or long_press actions. Such identifiers can be retrieved by determining which accessibility node encompasses the click location. When multiple nodes satisfy this criterion, the node with the smallest area is selected. However, we discard any step with an element-based action (click, long_press) that does not have a UI element associated. This is due to either a touch on an empty area or a target UI element missing from the accessibility tree. These discarded steps are still considered in the action history to support prediction of later steps that reference previous actions or elements. (iii) Finally, we discard (only from SeqIO LL tasks) steps that are missing a low-level instruction which annotators may have forgotten to enter. An episode is dropped only if all of its steps are discarded. Table 7 shows the statistics of the SeqIO tasks that are the results of this processing.
Table 7: Statistics of the SeqIO tasks generated from AndroidControl. Note that the number of steps here reported is different than what reported in Table 3 because it accounts for terminate and type actions which are not present in the original dataset (see C.2).
@@ D.2 LLM prompts for zero-shot experiments
For an LLM-based UI control agent, its prompt describes what the agent is expected to do, the action space and the expected format of actions, the task instruction, the current screen and application, and the history of previously executed actions and their outcome.
@@ D.3 Action matching
In computing step-wise accuracy, we consider an action correctly predicted if it matches the ground truth action exactly (i.e., action type and arguments are identical) or if it aligns with it as described below.
@@ E.2 Training with different levels of instructions
We observe the screen representations provided as input to the UI control agent are critical to its success. Table 13 compares the step accuracy of various models on Google’s first party apps and apps developed by third party developers (3rd-party). It is evident the gap in performance, especially in handling high-level instructions. All zero-shot methods perform significantly better on first-party apps than on third-party apps. This is evident for fine-tuned models, especially on high-level instructions where the most performance model (LT-all) achieves 82.5% step accuracy on first-party apps and only 58.7% accuracy on third-party apps. When the input is made of low-level instructions the gap is smaller and in some cases (few-shots) the performance on third-party apps is higher. This shows how for tasks that require stronger reasoning capabilities accurate screen representations are particularly critical.